OpenStudy (anonymous):

what is the area of rhombus with diagonals of 10ft and 14ft? PLEASE HELP.

OpenStudy (btaylor):

the area is half the product of the diagonals. So 1/2 (10 s 14) = 1/2 ( 140) = 70

OpenStudy (anonymous):

Yay. Thank you. :)